If a car is driving at 25 mph, how long will it take that car to travel 320ft?
Answer by stanbon(75887)

If a car is driving at 25 mph, how long will it take that car to travel 320ft?
--------
time = distance/rate
----
320ft = 320/5280 = 0.06061 miles
time = 320/25 [ft/(mph)] = 0.06061 miles/25mph = 0.0024 hr = 8.73 seconds
---------------------
Cheers,
Stan H.